### v1

@@ Abstract
We report an implementation of the McMurchie-Davidson evaluation scheme for 1- and 2-particle Gaussian AO integrals designed for efficient execution on modern central processing units (CPUs) with Single Instruction Multiple Data (SIMD) instruction sets. Like in our recent MD implementation for graphical processing units (GPUs) [J. Chem. Phys. 160, 244109 (2024)], variable-sized batches of shellsets of integrals are evaluated at a time. By optimizing for the floating point instruction throughput rather than minimizing the number of operations, this approach achieves up to 50% of the theoretical hardware peak FP64 performance for many common SIMD-equipped platforms (AVX2, AVX512, NEON), which translates to speedups of up to 30 over the state-of-the-art one-shellset-at-a-time implementation of Obara-Saika-type schemes in Libint for a variety of primitive and contracted integrals. As with our previous work, we rely on the standard C++ programming language — such as the std::simd standard library feature to be included in the 2026 ISO C++ standard — without any explicit code generation to keep the code base small and portable. The implementation is part of the open source LibintX library freely available at https://github.com/ValeevGroup/libintx.
@@ 1 Introduction
The key challenge of exploiting the wide SIMD vector units of modern CPUs is the lack of a regular source of data parallelism in the Gaussian integral kernels. Integral engines such as Libcint that use the Rys quadrature scheme can exploit data parallelism over the quadrature roots while still evaluating one shellset at a time.6 A recent report suggested that such strategy is effective only for short vectors (SSE, not AVX) due to the inability to fill the entire SIMD vector with work except for integral classes with high angular momenta.7 For auxiliary-function evaluation schemes such as Obara-Saika and McMurchie-Davidson, vectorization over individual integrals in a shell set is not possible due to the irregular arithmetic structure of such compute schemes. Furthermore, the contraction degrees of Gaussian AOs in most practical basis sets are usually too low to be able to populate a whole SIMD vector with primitive shellsets of a single contracted shellset. Therefore, one must consider evaluation of multiple shellsets at a time. Pritchard and Chow explored vectorization over the primitives of the vertical recurrence part of the Obara-Saika-Head-Gordon-Pople (OSHGP) scheme8, 9 when computing multiple shellsets at a time.10 On an AVX core their implementation achieved speedups of $\sim 2$ over the scalar code and over Libint for the augmented correlation-consistent triple-zeta basis (with medium contraction degree) and higher speedups for deeply-constracted atomic natural orbital triple-zeta basis; unfortunately no results were reported for basis sets with low contraction degrees, such as the Turbomole def2 basis set family.11
@@ 2.1 McMurchie-Davidson Scheme
denotes an uncontracted primitive Cartesian Gaussian with exponent $\zeta_{a}\in\mathbb{R}^{+}$ and non-negative Cartesian “quanta” $\mathbf{a}\equiv\{a_{x},a_{y},a_{z}\},a_{i}\in\mathbb{Z}$ centered at $\mathbf{A}\equiv\{A_{x},A_{y},A_{z}\}$, where $\mathbf{r}_{A}\equiv\{x_{A},y_{A},z_{A}\},x_{A}\equiv x-A_{x}$, etc. The sum of Cartesian quanta, $l_{\mathbf{a}}\equiv a_{x}+a_{y}+a_{z}\geq 0$, is colloquially referred to as the “angular momentum” of a Gaussian.
@@
needed for various approximations to the 2-particle Coulomb interaction (density fitting13, 14, 15 and others)16, 17 and for computing empirical atomic potentials in the context of guess orbital construction.18 The 1- and 2-particle integrals will also be generically denoted as $[\mathrm{bra}]_{\hat{O}}$ and $[\mathrm{bra}|\mathrm{ket}]$, respectively.
@@
[…]
@@
[…]
@@
with Hermite Gaussian exponent and origin for the 1- and 2-center bra given by $\zeta_{p}\equiv\zeta_{a}$, $\mathbf{P}\equiv\mathbf{A}$ and $\zeta_{p}\equiv\zeta_{a}+\zeta_{b}$, $\mathbf{P}\equiv\frac{\zeta_{a}\mathbf{A}+\zeta_{b}\mathbf{B}}{\zeta_{a}+\zeta_{b}}$, respectively. Hermite Gaussian used to expand the 2-center ket will be denoted by $\tilde{\mathbf{q}}$, with parameters $\zeta_{q}\equiv\zeta_{c}+\zeta_{d}$, $\mathbf{Q}\equiv\frac{\zeta_{c}\mathbf{C}+\zeta_{d}\mathbf{D}}{\zeta_{c}+\zeta_{d}}$.
@@
$\left(E_{i}\right)_{0}^{0}$ $=$ $1,$ (12)
$E_{a_{i}}^{\tilde{p}_{i}}$ $=$ $0\qquad(\tilde{p}_{i}\notin[0,a_{i}]\vee\mod(\tilde{p}_{i}+a_{i},2)=1).$ (13)
Similarly, recurrences Eqs. 10 and 11 are bootstrapped by
@@
[…]
@@
$H_{\mathbf{a}}^{\mathbf{\tilde{p}}}=$ $\sum_{\mathbf{p}}C_{l_{a}m_{a}}^{\mathbf{a}}E^{\mathbf{\tilde{p}}}_{\mathbf{a}},$ (18)
@@
Shells are groups of solid harmonic and Hermite Gaussians that share exponents, origin, contraction coefficients, and orbital quantum numbers. There are $12l+$ 1 and $(l+1)(l+2)(l+3)/6$ solid harmonic and Hermite Gaussians in a shell of angular momentum $l$, respectively. Shells of angular momenta {0, 1, 2, 3, 4, 5, 6} will be denoted by {s, p, d, f, g, h, i}, or by their angular momentum itself. E.g., [21||||43] and [dp||||gf] will denote a class of integrals representing any shellset of 4-center integrals with d, p, g, and f shells on each respective center.
@@
$[\mathbf{\tilde{p}}|\mathbf{\tilde{q}}]\equiv$ $\,\iint\frac{\Lambda_{\mathbf{\tilde{p}}}(\mathbf{r}_{1})\Lambda_{\mathbf{\tilde{q}}}(\mathbf{r}_{2})}{|\mathbf{r}_{1}-\mathbf{r}_{2}|}\,\mathrm{d}\mathbf{r}_{1}\,\mathrm{d}\mathbf{r}_{2}.$ (21)
Both can be obtained from the auxiliary integrals,
@@
$[\mathbf{0}]_{\mathbf{I}}^{(m)}$ and $[\mathbf{0}]^{(m)}$ are related to the Boys function $F_{m}(x)$ (or similar quantities for non-Coulomb integrals21):
@@
starting from $[\mathbf{0}]_{\mathbf{I}}^{(m)}$ and $[\mathbf{0}]^{(m)}$, respectively.
@@
The complete sequences of transformations involved in the MD scheme for evaluation of primitive Coulomb 1-particle and 2-particle integrals are
@@ 2.2.2 Early contraction of the $E_{\mathbf{a}\mathbf{b}}^{\mathbf{a}+\mathbf{b}}$ contribution
[…]
@@ 2.2.3 Coulomb 2-particle 3-center integrals
Note that the required range of $m$ in $[0]^{(m)}$ matches the range required by the Ahlrichs’ improvement of the OS scheme for 3-center 2-particle integrals over solid harmonic Gaussians.23 The match is not coincidental as both Ahlrichs’ and our optimizations omit the intermediate Cartesian/Hermite integrals that produce contributions to the target $[\mathbf{a}|\mathbf{cd}]$ integral with the overly slow multipolar decay, and hence such contributions are filtered out by the transformation to the solid harmonic basis.
This optimization reduces the entire hermite-to-AO transformation of the auxiliary bra to just that of the terminal term, with associated operations and memory savings. Note that the partial early contraction optimization described above for the 2-center ket could be used to calculate the entire contracted bra in this case. However, since the density fitting AOs are usually uncontracted, this will result in rather modest actual improvement. This idea has been described by Knizia et al.24 for 2-center 2-particle integrals over solid harmonic Gaussians but generalization to the 3-center case was not reported.
@@ 3.1 Basics
The new SIMD-capable MD implementation for standard 1-particle and up to 4-center 2-particle integrals was implemented in the LibintX library, freely available at github.com:ValeevGroup/LibintX (branch feature/md/simd) under the Lesser GNU Public License (version 3). As in our previous2, 1, 25, 3 work the library is written in standard C++; there is no custom code generator. Modern metaprogramming facilities of C++ allow to reach high performance while keeping the source code compact (fewer than 1,500 lines in total). Similar to the GPU implementation3 the integral kernel compilation is broken into multiple object targets. Special attention was paid to ensure that time and memory requirements for compilation remain reasonable even for high angular momenta. For example, compilation of the library configured to support {3,4}-center integrals with $l=6$ AOs takes {5,15} minutes using the LLVM Clang C++ compiler on four i7 cores, using <<<1GB of memory per translation unit.
@@
The CPU and GPU integral engines share the same interface via virtual base class. Both produce integrals in the same $[ij,ab,cd,kl]$ column-major layout (with $ij$ and $kl$ the bra and ket shell-pair indices and $ab$ and $cd$ the bra and ket AO pairs; see Ref. 3 for the layout discussion of the GPU implementation). Internally, integrals are computed in small batches (multiples of the SIMD vector size) and the CPU side has an interface to “consume” (process) those small batches on the fly rather than populating the entire integral array.
@@ 3.2.1 Boys functions
In evaluation of the Boys function for SIMD argument types both asymptotic and interpolation branches need to be evaluated with “invalid” SIMD lanes masked with zeroes. With AVX2 instructions it is possible to implement faster interpolation (by roughly 2×2$\times$2 ×) by rearranging the 7th order Chebyshev interpolation into two parts and using vector permutation intrinsics. Such optimization is not possible with 512-bit instructions as the vector size is equal to the number of interpolation coefficients. In fact, this AVX2 implementation is faster than the generic std::simd AVX512 version; therefore, the AVX2 implementation is used regardless of the vector width. This is the only place where architecture-specific code is used; interested readers are referred to the source code.
@@
For 3-center 2-particle integrals, the 1-center bra $[X|q]$ is transformed first, followed by ket transformation to $[X|cd]$. If ket fits into registers (for example, ||||dp] fits into 16 registers), a small inline matrix kernel is used, the BLAS GEMM is used otherwise. If BLAS GEMM is used, $X$ centers are batched together for better performance. The bra early contraction optimization (Section 2.2.3) is only used with BLAS GEMM at the moment as it inhibits the compiler from some inline code optimizations. The partial ket early contraction is not used as it did not produce much improvement.
The 4-center integral is evaluated as $[r]\to[p|cd]\to[ab|cd]$, with vectorization over bra as in the 3-center case. As with the 3-center code, larger Hermite-to-AO transforms are handled with BLAS GEMM, smaller with inline matrix kernel. The partial early contraction optimization (Section 2.2.2) is always used. Scalar bra transformation can be done with GEMM as well, but with SIMD leading dimension is now interleaved vector elements, requiring explicit transpose steps to use GEMM (either before or after). This also prevents us from using small matrix libraries like libxsmm. Instead, SIMD bra transformation uses another bespoke inline matrix kernel. Bra will tend to dominate where the ket is uncontracted and is much smaller, e.g., (ii||||pp), but those cases are less common. Notice that in such cases BLAS GEMM is (expectedly) faster than the bespoke kernel.

### v2

@@ Abstract
We report an implementation of the McMurchie-Davidson evaluation scheme for 1- and 2-particle Gaussian AO integrals designed for processors with Single Instruction Multiple Data (SIMD) instruction sets. Like in our recent MD implementation for graphical processing units (GPUs) [J. Chem. Phys. 160, 244109 (2024)], variable-sized batches of shellsets of integrals are evaluated at a time. By optimizing for the floating point instruction throughput rather than minimizing the number of operations, this approach achieves up to 50% of the theoretical hardware peak FP64 performance for many common SIMD-equipped platforms (AVX2, AVX512, NEON), which translates to speedups of up to 30 over the state-of-the-art one-shellset-at-a-time implementation of Obara-Saika-type schemes in Libint for a variety of primitive and contracted integrals. As with our previous work, we rely on the standard C++ programming language — such as the std::simd standard library feature to be included in the 2026 ISO C++ standard — without any explicit code generation to keep the code base small and portable. The implementation is part of the open source LibintX library freely available at https://github.com/ValeevGroup/libintx.
@@ 1 Introduction
The key challenge of exploiting the wide SIMD vector units of modern CPUs is the lack of a regular source of data parallelism in the Gaussian integral kernels. Integral engines such as Libcint that use the Rys quadrature scheme can exploit data parallelism over the quadrature roots while still evaluating one shellset at a time.6 A recent report suggested that such strategy is effective only for short vectors (SSE, not AVX) due to the inability to fill the entire SIMD vector with work except for integral classes with high angular momenta.7 For auxiliary-function evaluation schemes such as Obara-Saika and McMurchie-Davidson, vectorization over individual integrals in a shell set is not possible due to the irregular arithmetic structure of such compute schemes. Furthermore, the contraction degrees of Gaussian AOs in most practical basis sets are usually too low to be able to populate a whole SIMD vector with primitive shellsets of a single contracted shellset. Therefore, one must consider evaluation of multiple shellsets at a time. Pritchard and Chow explored vectorization over the primitives of the vertical recurrence part of the Obara-Saika-Head-Gordon-Pople (OSHGP) scheme8, 9 when computing multiple shellsets at a time.10 On an AVX core their implementation achieved speedups of $\sim 2$ over the scalar code and over Libint for the augmented correlation-consistent triple-zeta basis (with medium contraction degree) and higher speedups for deeply contracted atomic natural orbital triple-zeta basis; unfortunately no results were reported for basis sets with low contraction degrees, such as the Turbomole def2 basis set family.11
@@ 2.1 McMurchie-Davidson Scheme
denotes an uncontracted primitive Cartesian Gaussian with exponent $\zeta_{a}\in\mathbb{R}^{+}$ and non-negative Cartesian “quanta” $\mathbf{a}\equiv\{a_{x},a_{y},a_{z}\},a_{i}\in\mathbb{Z}$ centered at $\mathbf{A}\equiv\{A_{x},A_{y},A_{z}\},A_{i}\in\mathbb{R}$, where $\mathbf{r}_{A}\equiv\{x_{A},y_{A},z_{A}\},x_{A}\equiv x-A_{x}$, etc. The sum of Cartesian quanta, $l_{\mathbf{a}}\equiv a_{x}+a_{y}+a_{z}\geq 0$, is colloquially referred to as the “angular momentum” of a Gaussian.
@@
needed for various approximations to the 2-particle Coulomb interaction (density fitting13, 14, 15 and others)16, 17 and for computing empirical atomic potentials in the context of guess orbital construction.18 The 1- and 2-particle integrals will also be generically denoted as $\mathrm{bra}]_{\hat{O}}$ and $\mathrm{bra}|\mathrm{ket}]$, respectively.
@@
with $K_{a}\geq 1$ denoted by contraction degree. In the standard McMurchie-Davidson (MD) scheme12 used here the integrals are computed in the primitive basis and accumulated into the contracted integrals; this is different from the Head-Gordon-Pople refinement9 of the Obara-Saika scheme8 implemented in Libint where nontrivial work is performed in the contracted basis. The contracted integrals will be denoted by the use of parentheses instead of square brackets. The total contraction degree of an integral is the product of the contraction degrees of the constituent AOs. For 2-particle integral $\text{bra}|\text{ket}]$ the contraction degrees of the bra and ket contraction degrees will be denoted as $K_{\text{bra}}$ and $K_{\text{ket}}$, respectively; clearly, $K=K_{\text{bra}}K_{\text{ket}}$.
@@
$\phi_{\mathbf{a}}({\bf r})=$ $\sum_{\tilde{a}_{x}=0}^{\tilde{p}_{x}\leq a_{x}}E_{a_{x}}^{\tilde{p}_{x}}\sum_{\tilde{p}_{y}=0}^{\tilde{p}_{y}\leq a_{y}}E_{a_{y}}^{\tilde{p}_{y}}\sum_{\tilde{p}_{z}=0}^{\tilde{p}_{z}\leq a_{z}}E_{a_{z}}^{\tilde{p}_{z}}\Lambda_{\mathbf{\tilde{p}}}(\mathbf{r})\equiv\sum_{\mathbf{\tilde{p}}}E_{\mathbf{a}}^{\mathbf{\tilde{p}}}\Lambda_{\mathbf{\tilde{p}}}(\mathbf{r}),$ (7)
@@
with Hermite Gaussian exponent and origin for the 1- and 2-center bra given by $\zeta_{p}\equiv\zeta_{a}$, $\mathbf{P}\equiv\mathbf{A}$ and $\zeta_{p}\equiv\zeta_{a}+\zeta_{b}$, $\mathbf{P}\equiv\frac{\zeta_{a}\mathbf{A}+\zeta_{b}\mathbf{B}}{\zeta_{a}+\zeta_{b}}$, respectively. Hermite Gaussian used to expand the 2-center ket will be denoted by $\tilde{\mathbf{q}}$, with parameters $\zeta_{q}\equiv\zeta_{c}+\zeta_{d}$, $\mathbf{Q}\equiv\frac{\zeta_{c}\mathbf{C}+\zeta_{d}\mathbf{D}}{\zeta_{c}+\zeta_{d}}$. The tilde over Cartesian quanta, such as $\tilde{\mathbf{a}}\equiv\{\tilde{a}_{x},\tilde{a}_{y},\tilde{a}_{y}\},\tilde{a}_{i}\in\mathbb{Z}$, will be used instead of the tilde-free counterparts $\tilde{\mathbf{a}}$ to distinguish Hermite Gaussians from their Cartesian counterparts; this will be especially useful in quantities involving both types of Gaussians.
@@
$\left(E_{i}\right)_{0}^{0}=1$ (12)
$E_{a_{i}}^{\tilde{p}_{i}}=0$ $,\quad\tilde{p}_{i}<0\textrm{ or }\tilde{p}_{i}>a_{i}\text{ or }(\tilde{p}_{i}+a_{i})\bmod{2}\neq 0$ (13)
Similarly, recurrences Eqs. 10 and $1$ are bootstrapped by
@@
$E_{a_{i}b_{i}}^{\tilde{p}_{i}}=$  $\,0,\quad\tilde{p}_{i}<0\textrm{ or }\tilde{p}_{i}>a_{i}+b_{i}\text{ or }(\tilde{p}_{i}+a_{i}+b_{i})\bmod{2}\neq 0$ (15)
@@
$H_{\mathbf{a}}^{\mathbf{\tilde{p}}}$= $\sum_{\mathbf{\tilde{p}}}C_{l_{a}m_{a}}^{\mathbf{a}}E^{\mathbf{\tilde{p}}}_{\mathbf{a}}$, (18)
@@
Shells are groups of solid harmonic and Hermite Gaussians that share exponents, origin, contraction coefficients, and orbital quantum numbers. There are $2l+1$ and $(l+1)(l+2)(l+3)/6$ solid harmonic and Hermite Gaussians in a shell of angular momentum $l$, respectively. Shells of angular momenta {0, 1, 2, 3, 4, 5, 6} will be denoted by {s, p, d, f, g, h, i}, or by their angular momentum itself. E.g., [21||43] and [dp||gf] will denote a class of integrals representing any shellset of 4-center integrals with d, p, g, and f shells on each respective center.
@@
$[\mathbf{\tilde{p}}|\mathbf{\tilde{q}}$$\equiv$ $\,\iint\frac{\Lambda_{\mathbf{\tilde{p}}}(\mathbf{r}_{1})\Lambda_{\mathbf{\tilde{q}}}(\mathbf{r}_{2})}{|\mathbf{r}_{1}-\mathbf{r}_{2}|}\,\mathrm{d}\mathbf{r}_{1}\,\mathrm{d}\mathbf{r}_{2}.$ (21)
Both can be obtained from the auxiliary integrals, [^1] Note that $\mathbf{r}$ refers to Cartesian coordinates of particles whereas $\tilde{\mathbf{r}}$ denotes Cartesian quanta of Hermite Gaussians.
@@
$\mathbf{0}]_{\mathbf{I}}^{(m)}$ and $\mathbf{0}]^{(m)}$ are related to the Boys function $F_{m}(x)$ (or similar quantities for non-Coulomb integrals21):
@@
starting from $\mathbf{0}]_{\mathbf{I}}^{(m)}$ and $\mathbf{0}]^{(m)}$, respectively.
@@
The complete sequences of transformations involved in the MD scheme for the evaluation of primitive Coulomb 1-particle and 2-particle integrals are
@@ 2.2.2 Early contraction of the $E_{\mathbf{a}\mathbf{b}}^{\mathbf{a}+\mathbf{b}}$ contribution
where $\sum_{\mathbf{cd}}C^{\mathbf{cd}}_{\mathrm{ket}}$ denotes the Cartesian-to-solid-harmonics transformation of the ket. Section 2.2.2 was previously exploited in our MD GPU integral engine3 to reduce the memory footprint of the auxiliary and Hermite integrals and the number of operations. This is especially significant for lower angular momenta. For example, for $l_{\text{ket}}=2$ (such as integrals with $|\text{pp}]$ or $|\text{ds}]$ kets) this technique allows to compute directly the contributions to Section 2.2.2 from 6 Hermite Gaussians with $l_{\tilde{\bf q}}=2$ which account for 60% of the total of 10 Hermite Gaussians in the d Hermite shell.
@@ 2.2.3 Coulomb 2-particle 3-center integrals
Note that the required range of $m$ in $0]^{(m)}$ matches the range required by the Ahlrichs’ improvement of the OS scheme for 3-center 2-particle integrals over solid harmonic Gaussians.23 The match is not coincidental as both Ahlrichs’ and our optimizations omit the intermediate Cartesian/Hermite integrals that produce contributions to the target $\mathbf{a}|\mathbf{cd}]$ integral with the overly slow multipolar decay, and hence such contributions are filtered out by the transformation to the solid harmonic basis.
This optimization reduces the entire hermite-to-AO transformation of the auxiliary bra to just that of the terminal term. Moreover if bra is contracted, early bra contraction reduces to contracting $\mathbf{0}]^{(m)}$ integrals rather than 1-index $\tilde{\mathbf{p}}]^{(0)}$ integrals as in Eq. 40. However, since the density fitting AOs are usually uncontracted, this will result in rather modest actual improvement. This idea has been described by Knizia et al.24 for 2-center 2-particle integrals over solid harmonic Gaussians but generalization to the 3-center case was not reported.
@@ 3.1 Basics
The new SIMD-capable MD implementation for standard 1-particle and up to 4-center 2-particle integrals was implemented in the LibintX library, freely available at github.com:ValeevGroup/LibintX (branch feature/md/simd) under the Lesser GNU Public License (version 3). As in our previous2, 1, 25, 3 work the library is written in standard C++; there is no custom code generator. Modern metaprogramming facilities of C++ allow to reach high performance while keeping the source code compact (fewer than 1,500 lines in total). Similar to the GPU implementation3 the integral kernel compilation is broken into multiple object targets. Special attention was paid to ensure that time and memory requirements for compilation remain reasonable even for high angular momenta. For example, compilation of the library configured to support {3,4}-center integrals with $l=6$ AOs takes {5,15} minutes using the LLVM Clang C++ compiler on four i7 cores, using <<1GB of memory per translation unit.
@@
The CPU and GPU integral engines share the same interface via virtual base class. Both produce integrals in the same $[ij,ab,cd,kl]$ column-major layout (with $ij$ and $kl$ the bra and ket shell-pair indices and $ab$ and $cd$ the bra and ket AO pairs; see Ref. 3 for the layout discussion of the GPU implementation). Internally, integrals are computed in small batches (multiples of the SIMD vector size) and the CPU side has an interface to “consume” (process) those small batches on the fly rather than populating the entire integral array. Screening is likewise done on a batch level with assumption that bra (and to lesser effect ket) indices are sorted in (any) order that has blocek sparsity.
@@ 3.2.1 Boys functions
In evaluation of the Boys function for SIMD argument types both asymptotic and interpolation branches need to be evaluated with “invalid” SIMD lanes masked with zeroes. With AVX2 instructions it is possible to implement faster interpolation (by roughly $2\times$) by rearranging the 7th order Chebyshev interpolation into two parts and using vector permutation intrinsics. Such optimization is not possible with 512-bit instructions as the vector size is equal to the number of interpolation coefficients. In fact, this AVX2 implementation is faster than the generic std::simd AVX512 version; therefore, the AVX2 implementation is used regardless of the vector width. This is the only place where architecture-specific code is used; interested readers are referred to the source code.
@@
For 3-center 2-particle integrals, the 1-center bra $[X|q]$ is transformed first, followed by ket transformation to $[X|cd]$. If ket fits into registers (for example, ||dp] fits into 16 registers), a small inline matrix kernel is used, the BLAS GEMM is used otherwise. If BLAS GEMM is used, $X$ centers are batched together for better performance. The bra early contraction optimization (Section 2.2.3) is only used with BLAS GEMM at the moment as it inhibits the compiler from some inline code optimizations. The partial ket early contraction is not used as it did not produce much improvement.
The 4-center integral is evaluated as $r]\to[p|cd]\to[ab|cd]$, with vectorization over bra as in the 3-center case. As with the 3-center code, larger Hermite-to-AO transforms are handled with BLAS GEMM, smaller with inline matrix kernel. The partial early contraction optimization (Section 2.2.2) is always used. Scalar bra transformation can be done with GEMM as well, but with SIMD leading dimension is now interleaved vector elements, requiring explicit transpose steps to use GEMM (either before or after). This also prevents us from using small matrix libraries like libxsmm. Instead, SIMD bra transformation uses another bespoke inline matrix kernel. Bra will tend to dominate where the ket is uncontracted and is much smaller, e.g., (ii||$p$), but those cases are less common. Notice that in such cases BLAS GEMM is (expectedly) faster than the bespoke kernel.